Unit B: 19 Aug 2011 activities

  • Present people:
    • Arcetri: Armando Riccardi, Marco Xompero
    • LBTO: Guido Brusa, Doug Miller, Michael Wagner, Richard Sosa

  • Activity log:
    • installation of 5mm shims on mirror cover of unit B to allow clearance between miniskirt and mirror cover even when hexapod is fully extracted (z=-12mm). The shims have been provided by ADS. The installation of the shims on mirror cover of unit A is still pending because mirror cover is currently on the telescope
    • shell box out of tent
    • cleaning of tent, mirror stand, mirror unit and tent floor. Fixed problems with clean room plastic wall stripes (excessive air leakage reducing overpressure)
    • unpacking and shell tray moved to the stand (in parallel to the cleaning activities)
    • miniskirt removal and cleaning of the reference body (in parallel to the unpacking activities):
      • Warning: 3 whitish spots found on the reference body out of capacitive sensor rings. The spots cannot be removed using Isopropanol or Acetone even rubbing with them. The spots seems to have no thickness (no step felt crossing the spot edges with a sharp piece of plastic, so spots have not been considered a problem. We went ahead.
    • stand with shell moved inside the tent (stans has wheels)
    • shell recentering and centering verification with central post
    • shell clocking check: clocking no more than 0.5 mm on the edge. No need to fix it
    • shell inspection: Guido found no chip, no cracks, no new relevant scratches
    • Installation of the reference signal distribution board on the shell:
      • electrical contact checked with a multimeter: OK

  • Guido and Richard left the mountain
  • We are currently on schedule

-- ArmandoRiccardi - 19 Aug 2011
Topic revision: r2 - 20 Aug 2011, ArmandoRiccardi
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback